# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v2

@@ Appendix B Proofs and Derivations
$\text{var}(\hat{\tau}^{(g)})$ =$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\right\}+\left\{\tilde{\beta}_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\right\}+\frac{2}{n_{g}}\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}$
Defining $\beta_{g}(z):=n_{g}\tilde{\beta}_{g}(z)$ (i.e., the reciprocal of the proportion of units in group $g$ who are assigned $Z_{i}=z$):
$=\frac{\beta_{g}(1)}{n_{g}}\text{var}_{g}\left\{Y_{i}(1)\right\}+\frac{\beta_{g}(0)}{n_{g}}\text{var}_{g}\left\{Y_{i}(0)\right\}$
$~{}~{}~{}~{}~{}-\frac{1}{n_{g}}\left[\text{var}_{g}\left\{Y_{i}(1)\right\}+\text{var}_{g}\left\{Y_{i}(0)\right\}-2\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}\right]$
$=\frac{\beta_{g}(1)}{n_{g}}\text{var}_{g}\left\{Y_{i}(1)\right\}+\frac{\beta_{g}(0)}{n_{g}}\text{var}_{g}\left\{Y_{i}(0)\right\}-\frac{1}{n_{g}}\text{var}_{g}(\tau_{i})$
@@
    The proportion of units in each subgroup converges to a proportion $\pi_{g}^{*}\in(0,1)$ (i.e., $n_{g}/n\to\pi_{g}^{*}$, where $10<\pi_{g}^{*}<$ 1 for all $g\in\{1,\ldots,G\}$ and $\sum_{g=1}^{G}\pi_{g}^{*}=1$.
@@
We would like to show that as $n\to\infty$, $\left(1-\frac{n_{z}}{n}\right)\cdot\frac{1}{n_{z}}\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))\to 0$.
Claim: $\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))=\frac{n}{n-1}\frac{1-\pi_{g}}{\pi_{g}}$.
@@ B.5 Proof of Corollary
[…]
@@
[…]
@@
[…]
B.7 Proof of Lemmas
B.7.1 Proof of Lemma B.1
@@ B.7 Proof of Lemmas
We will construct an upper bound on the first term ()(*)( ∗ ):
@@
$=\Pr\left(\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\right.$
@@
$\leq\Pr\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.\right.$
@@
$\leq\frac{1}{\delta_{ba}}\mathbb{E}\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.\right.$
@@ Appendix C Method Implementation Details
To perform honest estimation of the subgroup ATEs, we set $\pi_{train}=0.70$ for splitting the data into a training and hold-out estimation set. To estimate the individual-level treatment effects $\hat{\tau}(X_{i})$, we used the out-of-bag sample estimates (without repeated cross-fitting) in the Distilled Causal Forest and Distilled BCF, and we used the repeated cross-fitting procedure with $R=50$ repeats in the Distilled Rboost. We also examined other choices of $R$ in Figure A8. From these simulations, we found that $R=50$ provided a good balance between the subgroup estimation accuracy and computational burden. However, more generally, the subgroup estimation results from Distilled Rboost were robust as long as $R$ was sufficiently large enough.
@@ D.2 Additional Simulation Settings
Figure A3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes in the CATE-only outcome model scenario (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) frequently yields the highest $F_{1}$ and number of true positives alongside the lowest number of false positives, threshold RMSEs, and subgroup ATE RMSE, demonstrating its effectiveness for accurate subgroup estimation. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
Figure A5: Performance of pruned and unpruned versions of CDT and causal trees for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). While causal tree without pruning performs substantially worse than pruned causal tree, CDT with various teacher models (e.g., Distilled Causal Forest and Distilled Rboost) performs similarly well with or without pruning. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
Figure A6: Performance of subgroup estimation methods with oracle pruning for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). CDT with various teacher models remains the most robust and accurate subgroup estimation method, as seen by its high $F_{1}$ and number of true positives alongside low number of false positives, threshold RMSEs, and subgroup ATE RMSE. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
In the CDT algorithm, we have been using $R=50$ repeated cross-fits to estimate the individual-level treatment effects. To investigate the sensitivity of this choice, we show in Figure A8 the subgroup estimation performance of Distilled Rboost for varying choices of $R$, ranging between $1$ and $100$. In general, the subgroup estimation performance improves as $R$ increases and tends to be relatively stable for $R>10$.
@@
Figure A9: Performance of CDT with CART versus Rulefit student models for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). The Rulefit student model tends to estimate unnecessarily complex subgroups, compared to CART. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.

### v3

@@ Appendix B Proofs and Derivations
$\text{var}(\hat{\tau}^{(g)}\mid\mathcal{D})$ $=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\mid\mathcal{D}\right\}+\left\{\tilde{\beta}_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\mid\mathcal{D}\right\}+\frac{2}{n_{g}}\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\mid\mathcal{D}\right\}.$
@@
Proof.
We will prove consistency, and then derive the asymptotic variance. To show consistency, we take the expectation of $\hat{\tau}^{(g)}$.
Let $\overline{K_{\tau}}=\max_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left\{Y_{i}(1)-Y_{i}(0)\right\}$:
$\leq\tau^{(g)}+\frac{1}{n_{g}}\sum_{i=1}^{n}\overline{K_{\tau}}\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert$
To derive the variance, we apply law of total variance:
$\text{var}(\hat{\tau}^{(g)})=\mathbb{E}\left[\text{var}(\hat{\tau}^{(g)}\mid\mathcal{D})\right]+\text{var}\left[\mathbb{E}(\hat{\tau}^{(g)}\mid\mathcal{D})\right]$
We then examine both components of the variance expression. First,
$\mathbb{E}$ $\left[\text{var}(\hat{\tau}^{(g)}\mid\mathcal{D})\right]$
$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\mathbb{E}\left[\text{var}_{g}\left\{Y_{i}(1)\mid\mathcal{D}\right\}\right]+\left\{\tilde{\beta}_{g}(0)-\frac{1}{n_{g}}\right\}\mathbb{E}\left[\text{var}_{g}\left\{Y_{i}(0)\mid\mathcal{D}\right\}\right]+\frac{2}{n_{g}}\mathbb{E}\left[\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\mid\mathcal{D}\right\}\right]$
$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\right\}+\left\{\tilde{\beta}_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\right\}+\frac{2}{n_{g}}\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}$
where $\text{var}_{g}(A_{i}):=\mathbb{E}\left[\text{var}_{g}(A_{i}\mid\mathcal{D})\right]$.
Second, We can re-write the variance of $\mathbb{E}(\hat{\tau}^{(g)}\mid\mathcal{D})$ as:
var $\left[\mathbb{E}(\hat{\tau}^{(g)}\mid\mathcal{D})\right]$
$=$ $\text{var}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)$
$=$ $\text{var}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}+\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)$
$=$ $\underbrace{\text{var}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)}_{(1)}+\underbrace{\text{var}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)}_{(2)}$
$+2\underbrace{\text{cov}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\},\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)}_{(3)}$
Term (1) can be re-written as:
var $\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)$
$=\frac{1}{n_{g}}\left(\text{var}_{g}\{Y_{i}(1)\}+\text{var}_{g}\{Y_{i}(0)\}-2\text{cov}_{g}\{Y_{i}(1),Y_{i}(0)\}\right)$
For term (2):
var $\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)$
$=\frac{1}{n_{g}^{2}}\text{var}\left(\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\tau_{i}\right)$
$\leq\frac{1}{n_{g}}\frac{n}{n_{g}}\overline{K_{\tau}}^{2}\text{var}(\hat{\mathcal{G}}_{g}(X_{i})),$ (16)
where we have assumed $n_{g}/n\to\pi^{*}_{g}$ where $0<\pi^{*}_{g}<1$ for all $g\in\{1,...,G\}$. Then, Equation (16) converges to zero by Theorem 3.1. For term (3):
$\mathbb{E}$ $\left\{\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)\cdot\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)\right\}$
$=\frac{1}{n_{g}^{2}}\mathbb{E}\left[\left(\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\tau_{i}\right)\left(\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\tau_{i}\right)\right]$
$=\frac{1}{n_{g}^{2}}\left(\sum_{i=1}^{n}\mathbb{E}\left[\mathcal{G}_{g}(X_{i})\cdot\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\tau_{i}\right]+\sum_{i\neq j}\mathbb{E}\left[\mathcal{G}_{g}(X_{i})\{\hat{\mathcal{G}}_{g}(X_{j})-\mathcal{G}_{g}(X_{j})\}\tau_{i}\tau_{j}\right]\right)$
$=\frac{1}{n_{g}^{2}}\left(\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\tau_{i}\right]+\sum_{i\neq j}\mathcal{G}_{g}(X_{i})\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{j})-\mathcal{G}_{g}(X_{j})\}\tau_{i}\tau_{j}\right]\right)$
$=\frac{1}{n_{g}^{2}}\left(\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\tau_{i}\right]+\sum_{i\neq j}\mathcal{G}_{g}(X_{i})\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{j})-\mathcal{G}_{g}(X_{j})\}\tau_{i}\tau_{j}\right]\right)$
$\leq\frac{1}{n_{g}^{2}}\left(\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\overline{K_{\tau}}\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\right]+\sum_{i\neq j}\mathcal{G}_{g}(X_{i})\cdot\overline{K_{\rho}}\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{j})-\mathcal{G}_{g}(X_{j})\}\right]\right),$
where, $\overline{K_{\rho}}:=\max_{i\neq j}\tau_{i}\cdot\tau_{j}$. Under Theorem 3.1, this term will converge in probability to zero. We can similarly show that the lower bound of this term will similarly converge in probability to zero:
$\mathbb{E}$ $\left\{\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)\cdot\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)\right\}$
$\geq\frac{1}{n_{g}^{2}}\left(\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\underline{K_{\tau}}\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\right]+\sum_{i\neq j}\mathcal{G}_{g}(X_{i})\cdot\underline{K_{\rho}}\cdot\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{j})-\mathcal{G}_{g}(X_{j})\}\right]\right),$
which will converge in probability to zero under Theorem 3.1. Then, by Squeeze theorem, we have shown:
$\mathbb{E}\left\{\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)\cdot\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)\right\}\stackrel{{\scriptstyle p}}{{\rightarrow}}0.$
Similarly, we can show:
$\mathbb{E}$ $\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\}\right)\cdot\mathbb{E}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)$
$=\tau^{(g)}\cdot\frac{1}{n_{g}}\sum_{i=1}^{n}\mathbb{E}\left[\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\tau_{i}\right].$
Then, again, using the boundedness of $\tau_{i}$ (i.e., $\underline{K_{\tau}}\leq\tau_{i}\leq\overline{K_{\tau}}$), we can apply Squeeze theorem to show:
$\stackrel{{\scriptstyle p}}{{\rightarrow}}0.$
As such, we have shown:
$\text{cov}\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\{Y_{i}(1)-Y_{i}(0)\},\frac{1}{n_{g}}\sum_{i=1}^{n}\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\}\cdot\{Y_{i}(1)-Y_{i}(0)\}\right)\stackrel{{\scriptstyle p}}{{\rightarrow}}0$
Combining the terms together, we have shown:
$\text{asyvar}(\hat{\tau}^{(g)})$ $=\tilde{\beta}_{g}(1)\text{var}_{g}\{Y_{i}(1)\}+\tilde{\beta}_{g}(0)\text{var}_{g}\{Y_{i}(0)\}.$
B.5 Proof of Theorem A.2
@@
    The proportion of units in each subgroup converges to a proportion $\pi_{g}^{*}\in(0,1)$ (i.e., $n_{g}/n\to\pi_{g}^{*}$, where $0<\pi_{g}^{*}<1$ for all $g\in\{1,\ldots,G\}$ and $\sum_{g=1}^{G}\pi_{g}^{*}=1$.
@@
We will show that as $n\to\infty$, $\left(1-\frac{n_{z}}{n}\right)\cdot\frac{1}{n_{z}}\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))\to 0$.
First, note:
$\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))=\frac{n}{n-1}\frac{1-\pi_{g}}{\pi_{g}}.$
@@
$\text{asyvar}(\hat{s}_{n})=\left(dF_{X}(s)\sigma^{2}\right)^{2}\cdot\frac{2^{-2/3}}{6\pi i}\left(\frac{1}{2}\frac{V}{dF_{X}(s)\sigma^{2}}\right)^{-4/3}\int_{-\infty}^{\infty}\frac{t}{Ai(it)^{2}}dt,$
@@
$g_{\gamma}(t)$ $=\left(\frac{2}{\gamma}\right)^{1/3}\frac{1}{2\pi i}\int_{c_{1}-i\infty}^{c_{1}+i\infty}\frac{\exp(-tu)}{\text{Ai}((2\gamma^{2})^{-1/3}u)}du,$ (17)
@@
$\text{var}(Z_{\gamma})=\frac{2^{-2/3}\gamma^{-4/3}}{6\pi i}\int_{-\infty}^{\infty}\frac{t}{\text{Ai}(it)^{2}}dt.$
@@
$\text{asyvar}(\hat{s}_{n})=\left(dF_{X}(s)\sigma^{2}\right)^{2}\cdot\frac{2^{-2/3}}{6\pi i}\left(\frac{1}{2}\frac{V}{dF_{X}(s)\sigma^{2}}\right)^{-4/3}\int_{-\infty}^{\infty}\frac{t}{Ai(it)^{2}}dt.$
B.8 Proof of Lemmas
B.8.1 Proof of Lemma B.1
@@ B.8 Proof of Lemmas
We will construct an upper bound on the first term ()(*):
@@
[…]
@@
[…]
@@
[…]
@@ Appendix C Method Implementation Details
To perform honest estimation of the subgroup ATEs, we set $\pi_{train}=0.70$ for splitting the data into a training and hold-out estimation set. To estimate the heterogeneous treatment effects $\hat{\tau}(X_{i})$, we used the out-of-bag sample estimates (without repeated cross-fitting) in the Distilled Causal Forest and Distilled BCF, and we used the repeated cross-fitting procedure with $R=50$ repeats in the Distilled Rboost. We also examined other choices of $R$ in Figure A8. From these simulations, we found that $R=50$ provided a good balance between the subgroup estimation accuracy and computational burden. However, more generally, the subgroup estimation results from Distilled Rboost were robust as long as $R$ was sufficiently large enough.
@@ D.2 Additional Simulation Settings
Figure A3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes in the CATE-only outcome model scenario (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) frequently yields the highest $F_{1}$ and number of true positives alongside the lowest number of false positives, threshold RMSEs, and subgroup ATE RMSE, demonstrating its effectiveness for accurate subgroup estimation. Results are averaged across 100 simulation replicates with ribbons denoting $\pm 1SE$.
@@
Figure A5: Performance of pruned and unpruned versions of CDT and causal trees for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). While causal tree without pruning performs substantially worse than pruned causal tree, CDT with various teacher models (e.g., Distilled Causal Forest and Distilled Rboost) performs similarly well with or without pruning. Results are averaged across 100 simulation replicates with ribbons denoting $\pm 1SE$.
@@
Figure A6: Performance of subgroup estimation methods with oracle pruning for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). CDT with various teacher models remains the most robust and accurate subgroup estimation method, as seen by its high $F_{1}$ and number of true positives alongside low number of false positives, threshold RMSEs, and subgroup ATE RMSE. Results are averaged across 100 simulation replicates with ribbons denoting $\pm 1SE$.
@@
In the CDT algorithm, we have been using $R=50$ repeated cross-fits to estimate the heterogeneous treatment effects. To investigate the sensitivity of this choice, we show in Figure A8 the subgroup estimation performance of Distilled Rboost for varying choices of $R$, ranging between $1$ and $100$. In general, the subgroup estimation performance improves as $R$ increases and tends to be relatively stable for $R>10$.
@@
Figure A9: Performance of CDT with CART versus Rulefit student models for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). The Rulefit student model tends to estimate unnecessarily complex subgroups, compared to CART. Results are averaged across 100 simulation replicates with ribbons denoting $\pm 1SE$.